OpenStudy (anonymous):

Solve 2x - y = -6 if the domain is {-2, -1, 0, 1}.

OpenStudy (shinalcantara):

Domain are the values of x.. To find the values for the range, all you have to do is plug in the values of x to the equation. ------------ 2x - y = -6 y = 2x + 6 (for convenience, use this in substituting the values of x) ------------ when x=-2 y = 2x + 6 y = 2(-2) + 6 y = -4 + 6 y = ? ------------ when x=-1 y = 2x + 6 y = 2(-1) + 6 y = -2 + 6 y = ? ------------- when x=0 y = 2x + 6 y = 2(0) + 6 y = 0 + 6 y = ? ------------- when x=1 y = 2x + 6 y = 2(1) + 6 y = 2 + 6 y = ?

OpenStudy (anonymous):

Thanks so much.